# Harry Yang1, Linfeng Zhang3, Qifeng Chen1🖂, Yue Ma1🖂

### v2

@@ 4.1 Experimental Setup
In our experiment, we use the open-sourced image generation model FLUX.1-[dev] [18] as our base, and conduct all experiments in PyTorch on an NVIDIA A100 GPU with 40GB memory. For hyperparameters, we set the timestep to 28, guidance scale to 2.0, and $k_{\text{front}}$ to 2. We also apply multi-ControlNet conditioning using Depth and Canny maps, with injection timing set in the normalized denoising interval $[0.1,0.7]$, and respective strengths of 2.5 and 3.5.
Method Image Quality Background Preservation Text Align
MasaCtrl [5] 5.83 23.54 125.36 20.84
PnpInversion [45] 6.11 24.77 102.91 19.23
Dit4Edit [10] 6.14 24.36 83.75 22.66
Flux-Kontext [19] 6.53 32.91 18.35 28.53
KV-Edit [61] 6.51 34.73 16.42 26.97
RF-Solver [51] 6.52 33.28 17.53 30.41
FlowEdit [17] 6.42 32.46 18.92 28.94
Ours 6.57 35.79 8.23 33.71
Table 1: Comparison with the state-of-the-art methods on ReShapeBench. We evaluate the assessment on the image quality, background preservation, and text align, all of them implemented using official open-source code.
@@
Table 2: Ablation study on different $k_{\text{front}}$. The results show that with the increase of $k_{\text{front}}$, the editing results exhibit a trend of first improving and then degrading, and the $k_{\text{front}}=2$ is the best. ControlNet modules were disabled in this experiment to isolate the effect of $k_{\text{front}}$.
4.2 ReShapeBench Construction
4.2.1 Limitations
While several benchmarks have been introduced to evaluate image editing methods [52, 15, 55], they present certain limitations for our use case. Notably, PIE-Bench [15], a prominent benchmark for prompt-driven editing with 700 images, has two key shortcomings. First, its concise prompts lack the detail required for fine-grained, shape-aware edits. Second, its broad scope—covering object replacement, stylization, and background changes—makes it less focused for the specific evaluation of shape transformation capabilities. To overcome these issues, we introduce ReShapeBench.
4.2.2 Image Collection
ReShapeBench comprises 120 newly collected images organized into two primary subsets and one evaluation set. The first subset contains 70 images from online sources, each featuring a single, prominent object with well-defined boundaries suitable for precise shape editing. The second subset consists of 50 images with multiple objects, designed to test a model’s ability to perform targeted, mask-free edits. Finally, a general evaluation set of 50 images was created by curating a selection from the first two subsets and adding high-quality examples from PIE-Bench. This set is designed to assess model generalization across diverse shape-editing scenarios. All images are standardized to a resolution of $512\times 512$.
4.2.3 Refined Text Prompts
To address the issue of overly simple prompts in existing benchmarks, every image in ReShapeBench is paired with detailed source and target prompts. Both prompt types adhere to a consistent four-sentence structure: the first sentence provides a general summary, the second describes the foreground object(s), the third details the background, and the fourth describes the overall scene. When creating edit prompts, only the relevant attributes (such as object identity or features) are modified. These prompts are manually crafted to specify significant shape transformations, such as changes to contours or proportions, that result in low spatial alignment with the source object. This level of detail enables accurate, mask-free object recognition. All prompts were initially generated using Qwen-2.5-VL [3] and subsequently validated by hand to ensure precision and contextual consistency. Further details are available in the Supplementary Materials.
4.3 Comparison with Baselines
4.3.1 Qualitative Comparison.
We compare Follow-Your-Shape with two categories of image editing methods: Diffusion-based and Flow-based methods. Diffusion-based baselines include PnPInversion [15], MasaCtrl [5], and Dit4Edit [10], which perform editing by modulating attention mechanisms and conditions throughout the diffusion process. Flow-based baselines include RF-Edit [51], FlowEdit [17], KV-Edit [61], and FluxKontext [19], all of which build upon the Rectified Flow framework for controllable generation. Specifically, FluxKontext is a newly proposed model that leverages context-token concatenation, enabling strong in-context editing capabilities. Fig. 4 presents qualitative results showcasing the superior shape-aware editing ability and background preservation of our method compared to these baselines. Diffusion-based methods usually struggle to preserve the background under structural edits (e.g., MasaCtrl and Dit4Edit for the ’lion-shaped latten’) and sometimes fail to perform high-magnitude shape transformations (e.g., PnPInversion for the ’car leaps through the air’). On the other hand, flow-based methods generally produce higher-quality images with better background preservation, but still suffer from issues such as detail jitter (e.g., Flux-Kontext for the ’car leaps through the air’, KV-Edit for the ’lion-shaped latten’), ghosting artifacts (e.g., KV-Edit for the ’dragonflies’), and failure to complete large-scale shape transformations in challenging cases (e.g., all baselines for the ’hat’). In contrast, Follow-Your-Shape excels at performing large-scale shape transformations while faithfully preserving non-target regions.
4.3.2 Quantitative comparison.
Following the same categorization, we conduct the quantitative comparison on ReShapeBench to evaluate the effectiveness of our method against both diffusion-based and flow-based baselines. To ensure fairness, we use identical source and target prompts and apply the same number of denoising timesteps across all methods. Specifically, as we’re following the RF-Solver implementation and utilizing its second-order solver, we double the number of timesteps for methods that do not employ a second-order scheme to ensure the same NFE (Number of Function Evaluations). As shown in Tab. 1, we select the PNSR [14] and LPIPS [56] to evaluate the background consistency, Aesthetic Scores computed via the LAION aesthetic predictor [43] for the image quality, and the CLIP [40] for the text alignment. The results demonstrate that our model outperforms the baseline models across all metrics. The proposed region-controlled editing strategy achieves better performance for the shape-aware editing tasks, while our designed edit mask $M_{S}$ preserves the background well.
Figure 5: The qualitative ablation study on $k_{\text{front}}$. As the number of stabilization steps increases, background is better preserved, but shape transformation becomes less effective.
Figure 6: Ablation on ControlNet conditioning timestep and condition strength. (a) Conditioning is applied within five subranges of the normalized denoising interval $[0,1]$. Injection at relatively early stages (e.g., $[0.1,0.3]$) yields the most stable results. (b) Condition strength of the depth and canny branches, denoted as $(\text{depth},\text{canny})$, is varied across different settings. Excessive strength can overly constrain the edit, while insufficient strength leads to weak guidance.
@@ 4.4 Ablation Study
We conduct ablation studies to assess two key components in our editing framework: Initial trajectory stabilization steps and the timing and strength of ControlNet conditioning.
4.4.1 Effectiveness of Initial Trajectory Stabilization
To assess the role of initial trajectory stabilization, we vary the number of stabilization steps $k_{\text{front}}$ from 0 to 4. As shown in Figure 5, a small $k_{\text{front}}$ leads to noticeable drift and structural deviation from the source image, while a large $k_{\text{front}}$ overly restricts the intended shape transformation. To ensure a controlled quantitative comparison, ControlNet modules were disabled to isolate the effect of $k_{\text{front}}$. Table 2 shows that as $k_{\text{front}}$ increases, background preservation improves but CLIP similarity consistently drops, indicating a trade-off where trajectory stabilization reduces editing flexibility. Setting $k_{\text{front}}$ to 2 provides the best performance for preserving layout structure and allowing flexible edits.
4.4.2 Effectiveness of ControlNet conditioning timestep and strength
To explore the effect of ControlNet conditioning timestep, we vary the injection interval within the normalized denoising range $[0,1]$. Figure 6 (a) shows that injecting at relatively early stages yields the best, as the latent features are less noisy and more receptive to structural guidance. We also vary the conditioning strength on both Canny and Depth guidance. As shown in Figure 6 (b), a moderate scale (e.g., (2.5, 3.5)) achieves the best balance between structure preservation and editability, while overly weak or strong signals lead to under- or over-constrained outputs.
@@ 6.1 Rectified Flow (RF)
starting from a sample $X_{1}\sim\mathcal{N}(0,I)$. Since the ODE does not admit a closed-form solution in general, sampling is performed via numerical integration over a discretized set of timesteps $\{t_{i}\}_{i=0}^{N}$. A common approach is to use first-order solvers such as Euler or Heun’s method to approximate the trajectory:
@@
However, these first-order solvers can suffer from numerical instability and truncation error, especially in high-dimensional generation tasks. Several recent works [23, 42, 26, 8, 51] have explored higher-order integration strategies or adaptive solvers to improve generation fidelity. Specifically, RF-Solver [51] introduces a second-order update derived from a Taylor expansion of the velocity field:
@@ 6.2 KV Injection
When adapted to vision models, KV injection serves as a mechanism to enforce structural consistency by injecting key and value features from the inverted source image into the self-attention layers of the generative model. This guides attention toward the original spatial layout, helping to constrain edits to the desired region. In architectures like U-Net [41] and DiT [38], KV injection is typically applied during the denoising phase, where reused attention features act as anchors that preserve the overall scene structure.
To reduce memory cost and avoid limiting foreground flexibility, recent work such as StableFlow [2] explores the vital layers within DiT crucial for image formation. Therefore, by only reusing KV pairs in a subset of layers, it can balance structural fidelity and editability while effectively reduce the memory usage.
@@ 7.1 Shape Transformation
While recent generative models exhibit strong general-purpose editing capabilities, the concept of shape transformation remains ill-defined in the literature. Existing approaches typically perform localized edits—such as detail refinement, scaling, or texture changes—often guided by external signals like masks or ControlNet images. Although these methods succeed in achieving visible modifications, they are not explicitly framed as shape transformations and lack a clear taxonomy for distinguishing what qualifies as a true change in shape. When constructing ReShapeBench, we need a clear definition and categorization of shape transformation to guide data curation and enable meaningful evaluation.
By analyzing a wide range of prompt editing pairs, we identify a representative class of transformations in which the object’s contour, geometry, and even semantic identity undergo significant changes, yet its role in the scene remains intact. As shown in Figure 7, typical examples include ’a swan’ transformed into ’a small wooden boat’, ’a flamingo’, etc., where the object’s outline or category changes entirely, but the spatial location, visual salience, and narrative function of the object are preserved. Based on these observations, we define shape transformation of two objects as structural change that satisfies four key criteria below:
  • Cross-contour: The object’s overall silhouette or boundary changes significantly, beyond local warping or resizing.
  • Cross-semantic: The transformation changes the object into one from a different semantic class, while preserving overall scene consistency.
  • Structural transition: The transformation reflects a global reconfiguration of the object’s visual form, requiring changes to multiple parts more than attributes (e.g., colors, textures).
  • Subject continuity: Despite changes in shape and semantics, the object should maintain its spatial role and salience within the scene, ensuring visual coherence and contextual consistency.
The most challenging part in shape transformation is it requires the model to reinterpret object geometry while maintaining consistency in background, perspective, and composition. Unlike existing editing benchmarks that encompass on diverse editing tasks, our benchmark emphasizes large-scale shape transformation under prompt guidance, without relying on masks or external conditioning.
@@ 7.2 Image and Prompt Examples
Following the criteria of shape transformation outlined in the previous section, each prompt pair in ReShapeBench is carefully constructed to reflect meaningful transformations. For each selected source image, we first brainstormed plausible shape transformation targets that satisfy the criteria, then selected the most representative transformation for each image to ensure diversity and task difficulty across the dataset.
Prior benchmarks such as PIEBench often rely on single-sentence prompts, which are concise but frequently under-specify spatial or structural details. To generate high-quality prompts for our benchmark, we utilized the Qwen-VL-Chat [3] model. After specifying the intended transformation for a given image, we provided both the source prompt and the editing instruction to Qwen-VL-Chat. Leveraging its multi-turn dialogue memory, the model was able to maintain consistent formatting and descriptive style across all prompt pairs, producing aligned edit prompt. Examples are illustrated in Figure 9. Complete image data are included in the data appendix.
@@ 7.3 Evaluation Metrics
Aesthetic Score (AS) is used to assess the perceptual quality of the generated image, reflecting how well the visual content adheres to natural image statistics. This metric is computed using the LAION aesthetic predictor [43], which applies a linear estimator on top of CLIP embeddings to predict aesthetic quality [^1] https://github.com/LAION-AI/aesthetic-predictor. This is especially relevant in shape transformation tasks where the output may differ substantially from the source. In particular, AS serves as a useful proxy for detecting unnatural boundaries or blending artifacts that may arise during large-scale shape transitions, helping to evaluate whether the new shape is integrated smoothly into the scene.
@@
For background preservation, we adopt two widely used metrics that capture different aspects of visual similarity. Peak Signal-to-Noise Ratio (PSNR) measures low-level pixel fidelity, while Learned Perceptual Image Patch Similarity (LPIPS) evaluates perceptual similarity based on deep feature representations. Since we do not assume access to ground-truth masks, and the edited shape can vary across models, we follow the principle of subject continuity to localize the edited region. We apply a fixed-size box centered on the subject to occlude the foreground and compute similarity metrics over the remaining background area. This heuristic enables consistent and fair evaluation of how well the unedited context is preserved across different editing results.
@@
Finally, to assess text-image alignment, we compute CLIP similarity between the generated image and the target prompt. This metric offers a widely adopted embedding-based measure of semantic consistency.
Together, these metrics provide a comprehensive evaluation of structural realism, background integrity, and semantic alignment in shape-aware image editing.
Figure 9: Comparisons between original prompts and our refined prompts
Figure 10: Additional Shape-Aware Editing Results

### v4

@@ 5.1 Experimental Setup
We use the open-source FLUX.1-[dev] model [24] as the base and run all experiments in PyTorch on an NVIDIA A100 (40 GB). We set the number of denoising steps to 14, guidance scale to 2.0, and $k_{\text{front}}$ to 2. We evaluate both ControlNet-free and ControlNet-enabled variants of our method. If enabled, we apply multi-ControlNet conditioning with depth and Canny branches over the normalized denoising interval $[0.1,0.3]$, with respective strengths 2.5 and 3.5. Unless otherwise stated, we keep the same inference scheduler and tokenizer as the official release and fix random seeds for reproducibility; additional implementation details and runtime/memory profiles are provided in Supplementary Material 9.
5.2 Comparison with Baselines
5.2.1 Qualitative Comparison
We compare Follow-Your-Shape with diffusion-based and flow-based methods. Diffusion-based baselines include MasaCtrl [6], PnPInversion [20], and Dit4Edit [15], which modulate attention and conditions during the diffusion process. Flow-based baselines include RF-Edit [73], FlowEdit [22], KV-Edit [95], FLUX.1 Fill [24], and FLUX.1 Kontext [23], which build on Rectified Flow. FLUX.1 Fill is designed for prompt-based masked image completion, while FLUX.1 Kontext leverages context-token concatenation for in-context editing. Figure 6 shows that Follow-Your-Shape achieves stronger shape-aware editing and background preservation. Diffusion-based methods tend to degrade the background under structural edits and may fail on high-magnitude shape changes, while flow-based methods produce higher-quality images but still exhibit detail jitter, ghosting, or incomplete transformations in difficult cases. Follow-Your-Shape performs large-scale shape transformations while preserving non-target regions.
5.2.2 Quantitative comparison
Datasets ReshapeBench PIE-Bench
    Methods Metrics Image Quality Background Preservation Text Align Image Quality Background Preservation Text Align
[…]
MasaCtrl [6] 5.83 23.54 125.36 20.84 5.61 21.58 130.71 19.53
PnPInversion [20] 6.11 24.77 102.91 19.23 5.94 22.69 108.43 24.62
Dit4Edit [15] 6.14 24.36 83.75 22.66 6.03 22.74 97.65 23.87
RF-Edit [73] 6.52 33.28 17.53 30.41 6.49 31.97 15.34 29.67
FlowEdit [22] 6.42 32.46 18.92 28.94 6.37 32.68 16.42 28.93
KV-Edit [95] 6.51 34.73 16.42 26.97 6.47 33.45 13.72 28.14
FLUX.1Fill [24] 6.32 31.57 19.04 28.75 6.33 32.76 17.43 26.59
FLUX.1Kontext [23] 6.53 32.91 18.35 28.53 6.47 34.91 14.62 28.79
Ours (w/o ControlNet) 6.52 34.85 9.04 32.97 6.49 35.62 9.74 32.47
Ours (Full Model) 6.57 35.79 8.23 33.71 6.55 36.02 8.34 33.51
Table 1: Quantitative comparison with state-of-the-art methods on ReShapeBench and PIE-Bench. Bold and underlined values denote the best and second-best results respectively.
We conduct quantitative evaluations on both ReShapeBench and PIE-Bench against diffusion- and flow-based baselines to assess both shape-aware editing and general editing performances. To ensure fairness, we use identical source and target prompts and the same number of denoising steps across methods. Because we follow RF-Solver with a second-order scheme, we double the number of steps for methods without a second-order update to match the number of function evaluations (NFE). We disable the ControlNet modules to isolate the effect of TDM-guided editing.
@@ 5.2.2 Quantitative comparison
Table 2: Ablation study on different $k_{\text{front}}$. The results show that with the increase of $k_{\text{front}}$, the editing results exhibit a trend of first improving and then degrading, and the $k_{\text{front}}=2$ is the best.
As shown in Table 1, we evaluate background consistency with PSNR [19] and LPIPS [87], image quality with the LAION Aesthetic Score [61], and text alignment with CLIP similarity [58]. Supplementary Material 8.3 demonstrates the implementation details, including the preprocessing and metric computation settings used for all methods. Our method outperforms all baselines across metrics. Additionally, without the ControlNet module does not lead to a significant degradation in editing performance, indicating the improvements of our model are independent of this module. The region-controlled editing strategy improves fine-grained shape-aware editing while the mask $M_{S}$ preserves background content.
@@ 5.3 Ablation Study
Initial trajectory stabilization and the timing and strength of ControlNet conditioning have the largest impact on editing performance. We therefore ablate these two components: the former regulates early trajectory stability, while the latter controls structural guidance during mid–late steps.
Figure 7: (i) Ablation study on $k_{\text{front}}$. (ii.a) ControlNet conditioning applied within five subranges of the normalized denoising interval $[0,1]$. (ii.b) ControlNet conditioning strength of the depth and canny branches, denoted as $(\text{depth},\text{canny})$.
Figure 8: Visualization of ablation on $k_{\text{front}}$. While the aesthetic score remains relatively stable, PSNR, LPIPS, and CLIP reveal a clear trade-off between editing strength and background preservation.
5.3.1 Initial Trajectory Stabilization
To assess the role of initial trajectory stabilization, we vary the number of stabilization steps $k_{\text{front}}$ from 0 to 4. As shown in Figure 7 (i), small $k_{\text{front}}$ leads to drift and structural deviation, while large $k_{\text{front}}$ restricts the intended shape change. Table 2 shows that larger $k_{\text{front}}$ improves background preservation but reduces CLIP similarity, indicating a trade-off between stability and editability. $k_{\text{front}}=2$ provides the best balance, yielding stable trajectories while maintaining sufficient freedom for large shape transitions.
5.3.2 ControlNet Conditioning Timestep and Strength
To explore the effect of ControlNet conditioning timestep, we vary the injection interval within the normalized denoising range $[0,1]$. Figure 7 (ii.a) shows that earlier injection yields better results, as latent features are less noisy and more receptive to structural guidance. We also vary the depth and Canny strengths. As shown in Figure 7 (ii.b), moderate values (e.g., (2.5, 3.5)) best balance structure preservation and editability, while overly weak or strong signals under- or over-constrain the edit. These results suggest that early, moderate guidance best stabilizes geometry without suppressing desired semantic changes in the edited regions.
@@ 7.1 Rectified Flow (RF)
starting from a sample $X_{1}\sim\mathcal{N}(0,I)$. Since a closed-form solution is not available in general, we perform numerical integration over a discretized set of timesteps $\{t_{i}\}_{i=0}^{N}$. A standard choice uses first-order solvers such as Euler or Heun’s method to approximate the trajectory:
@@
However, first-order solvers can suffer from numerical instability and truncation error, especially in high-dimensional generation tasks. Several recent works [34, 60, 36, 9, 73] explore higher-order integration strategies or adaptive solvers to improve generation fidelity. Specifically, RF-Solver [73] introduces a second-order update derived from a Taylor expansion of the velocity field:
@@ 7.2 KV Injection
When extended from language to vision models, KV reuse often generalizes beyond strict token caching. In U-Net based models [6, 52], a common practice is to reuse intermediate attention maps or inject features derived from the inverted source image into self-attention layers. This feature-level injection plays a role similar to KV caching in LLMs by enforcing spatial consistency and anchoring the generative process to the source structure. In DiT-based architectures [50], this idea extends to reusing value (V) matrices or full KV pairs, providing finer-grained control over how structural information is preserved during denoising.
To reduce memory cost and avoid limiting foreground flexibility, recent work such as StableFlow [2] explores the vital layers within DiT crucial for image formation. Therefore, by only reusing KV pairs in a subset of layers, it can balance structural fidelity and editability while effectively reducing memory usage.
@@ 8.1 Shape Transformation
While recent image editing models exhibit strong general-purpose editing capabilities, the concept of shape transformation remains ambiguous in the literature. In practice, object modifications usually include detail adjustments, color changes, or limited geometric variations, often relying on masks or ControlNet images; however, such operations cannot be explicitly framed as shape transformations. When constructing ReShapeBench, we need a clear definition and categorization of shape transformation to guide data curation and enable meaningful evaluation.
From a geometric perspective, shape transformation is a structural change beyond local affine operations such as scaling, rotation, or minor warping. It reconfigures the object’s global contour and part topology, and may involve a shift in semantic class. At the same time, the transformed object must remain spatially coherent in the scene, occupying a similar anchor position and continuing to serve as the subject in context. Guided by these principles, we propose four criteria—cross-contour, cross-semantic, structural transition, and subject continuity—that together capture the essential properties of shape transformation.
  • Cross-contour: The object’s boundary undergoes a substantial change, exceeding local warping or affine resizing. This captures large-scale alterations to the external shape.
  • Cross-semantic: The transformation shifts the object into a different semantic class, indicating a categorical rather than attributive change, while preserving overall scene coherence.
  • Structural transition: The internal part topology is reconfigured, requiring modifications across multiple components instead of only simple attributes such as color or texture.
  • Subject continuity: The transformed object retains its spatial anchor and role in the scene, remaining contextually consistent despite the change in shape and semantic.
Almost every editing case in the paper can be classified as shape transformation. Figure 9 provides an additional illustrative example. Note that we exclude posture or viewpoint changes (standing $\rightarrow$ sitting), as these involve articulation or perspective variation rather than structural transformation. The most challenging part in shape transformation is that it requires the model to localize and reinterpret object shape while maintaining consistency in background and composition. Unlike existing editing benchmarks that cover diverse editing tasks, our benchmark emphasizes large-scale shape transformation under prompt guidance, without relying on masks or external conditioning.
@@ 8.2 Image and Prompt Examples
As described in Section 4, the benchmark is constructed from collected images divided into single-object and multi-object categories. Specifically, single-object cases broadly cover four categories—nature, animals, indoor, and outdoor scenes; multi-object cases can also be categorized into indoor and outdoor scenes. Figures 13 and 14 provide representative samples. For each image, source and target prompts follow the four-sentence template, which is illustrated in Table 5, 6, and 7.
@@ 8.3 Evaluation Metrics
Aesthetic Score (AS) measures the perceptual quality of the generated image by indicating how well the visual content conforms to natural image statistics. We compute AS with the LAION aesthetic predictor [61], which applies a linear estimator on top of CLIP embeddings [^1] https://github.com/LAION-AI/aesthetic-predictor. AS helps detect unnatural boundaries or blending artifacts that may occur during large-scale shape transitions and indicates how well the new shape integrates into the scene.
@@
For background preservation, we adopt two widely used metrics that capture different aspects of visual similarity. Peak Signal-to-Noise Ratio (PSNR) measures low-level pixel fidelity, while Learned Perceptual Image Patch Similarity (LPIPS) evaluates perceptual similarity based on deep feature representations. Since we do not assume access to ground-truth masks and the edited shape can vary across models, we localize the edited region by applying a fixed-size box centered on the subject to occlude the foreground and compute similarity over the remaining background area. This heuristic enables fair evaluation of how well the unedited content is preserved.
@@
Finally, to assess text-image alignment, we compute CLIP similarity between the generated image and the target prompt as an embedding-based measure of semantic consistency.
9 Implementation Details and More Ablation
9.1 Pseudocode for Follow-Your-Shape
Our pseudocode implementation is shown in Algorithn 1.
Algorithm 1 Region-Controlled Editing
Input: Inference steps $T$, Predicted velocities $\{v_{\text{src}}(\mathbf{x}_{t}^{(i)}),v_{\text{tgt}}(\mathbf{z}_{t}^{(i)})\}_{i=1}^{T}$, Source inversion features $\{K^{\text{inv}}_{t},V^{\text{inv}}_{t}\}_{t=0}^{T}$, target prompt $\mathbf{c}_{\text{tgt}}$, schedule phase durations $\{k_{\text{front}},k_{\text{tail}}\}$
[…]
2:for $t=T$ down to 11 do
3:   if $t>T-k_{\text{front}}$ then
4:     $M_{S}\leftarrow\mathbf{0}$
5:   else if $T-k_{\text{front}}\geq t>k_{\text{tail}}$ then
6:     $\delta_{t}^{(i)}\leftarrow\left\|v_{\text{tgt}}(\mathbf{z}_{t}^{(i)},t)-v_{\text{src}}(\mathbf{x}_{t}^{(i)},t,\mathbf{c}_{\text{src}})\right\|_{2}$
7:     $\tilde{\delta}_{t}^{(i)}\leftarrow\dfrac{\delta_{t}^{(i)}-\min_{j}\delta_{t}^{(j)}}{\max_{j}\delta_{t}^{(j)}-\min_{j}\delta_{t}^{(j)}}$ $\triangleright$ TDM Computation
8:     $M_{S}\leftarrow\mathbf{1}$
9:     $N\leftarrow N\cup\{t\}$
10:   else
11:     $\hat{\delta}^{(i)}\leftarrow\sum_{t^{\prime}\in T}\dfrac{\exp(\tilde{\delta}_{t^{\prime}}^{(i)})}{\sum_{t^{\prime\prime}\in T}\exp(\tilde{\delta}_{t^{\prime\prime}}^{(i)})}\cdot\tilde{\delta}_{t^{\prime}}^{(i)}$
12:     $\tilde{M}_{S}\leftarrow\mathcal{G}_{\sigma}*\hat{\delta} \triangleright$ TDM Aggregation
13:     $\tau\leftarrow\displaystyle\arg\max_{\tau^{\prime}}\;\mathbb{P}[\tilde{M}_{S}\leq\tau^{\prime}]\,\mathbb{P}[\tilde{M}_{S}>\tau^{\prime}]\,\big(\mathbb{E}[\tilde{M}_{S}\mid\tilde{M}_{S}\leq\tau^{\prime}]-\mathbb{E}[\tilde{M}_{S}\mid\tilde{M}_{S}>\tau^{\prime}]\big)^{2}$
14:     $M_{S}\leftarrow\mathbf{1}[M_{S}>\tau]$
15:   end if
16:   $K^{*}\leftarrow M_{S}\odot K^{\text{tgt}}_{t}+(1-M_{S})\odot K^{\text{inv}}_{t}$
17:   $V^{*}\leftarrow M_{S}\odot V^{\text{tgt}}_{t}+(1-M_{S})\odot V^{\text{inv}}_{t}$
18:end for
19:return $K^{*},V^{*}$
9.2 Hyperparameters
Hyperparameter Value
General
Inference Step 15
Guidance 2
Model Specific
$k_{\text{front}}$ 2
$k_{\text{tail}}$ 3
Softmax scale (temperature) 5
Gaussian smoothing $\sigma$ 0.7
Injecting DiT block (start idx) 19
Table 3: Hyperparameters used in our experiments.
The average running time for one image (averaged across multiple trials) is approximately 65.3 seconds, corresponding to 28 NFEs.
All hyperparameters used in our experiments can be seen in Table 3. The table is divided into two groups: General and Model Specific hyperparameters. Since our method is training-free, the only general hyperparameters on FLUX are the inference step and the guidance scale. Note that because we adopt RF solver, and the solver skips the final timestep, the number of function evaluations (NFE) is $(15-1)\times 2=28$. We have already performed the ablation studies about $k_{\text{front}}$ and ControlNet parameters (Refer to Section 5.3). $k_{\text{tail}}$ controls the number of late timesteps where source features are injected; setting it to 2 or 3 ensures that the model has sufficient time to perform editing while still converging to the target distribution. The softmax scale regulates the sharpness of the aggregated TDMs, with a higher value producing a crisper mask, while Gaussian smoothing $\sigma$ removes spurious noise to ensure the continuity of the mask. The injecting DiT block indicates the starting index of feature injection, and we found that injecting from block 19 onward achieves the best trade-off between editing quality and memory efficiency. This hyperparameter is relatively flexible and can be slightly adjusted depending on the particular editing case. To ensure reproducibility, we also provide the exact hyperparameters used for the specific examples shown in the paper (Table 4).
Task $k_{\text{front}}$ $k_{\text{tail}}$ CN type CN timing CN strength
parrot $\rightarrow$ hat, coconut, guitar, LOVE, rabbit 2 3 None NA NA
boy $\rightarrow$ cat, Teddy, backpack, soccer, lizard 1 3 Depth & Canny [0.1, 0.3] [2.5, 3.5]
bird $\rightarrow$ dragonflies, caps, robots, pokers 2 3 None NA NA
swan $\rightarrow$ boat, flamingo, crocodile, turtle 3 3 Depth [0.1, 0.3] 0.6
leaf latte $\rightarrow$ lion, horse $\rightarrow$ dragon, cocktail $\rightarrow$ ball 2 3 None NA NA
Table 4: Hyperparameters used for the results displayed in the paper.
9.3 Running Time and Memory Usage
Running Time Analysis.
The method requires an additional diffusion pass each step to compute the second-order prediction, resulting in a total of 28 NFEs. As introduced in Section 5.1, we conduct our experiment on an NVIDIA A100 (40 GB), and the average running time for one image (averaged across multiple trials) is approximately 65.3 seconds. The computational cost is comparable to existing methods such as FlowEdit and RF-Solver-Edit. Since our work focuses on controllability and structural fidelity rather than acceleration, no additional optimization loops are introduced. Speed optimization is orthogonal to our contributions.
Memory Usage.
During inference, the method stores a set of KV features and TDM maps on CPU memory, which amounts to approximately 12 GB in total. This cost is incurred only once during inversion and does not grow with the number of denoising steps. The GPU memory usage remains stable at around 25 GB throughout editing, comparable to existing flow-based editing pipelines. Since the method does not introduce any additional training procedures, this overhead is not a limiting factor in practice.
9.4 Post-hoc Analysis of Structural Localization Signals
To better understand whether the proposed Trajectory Divergence Map (TDM) provides a meaningful structural signal, we conduct a post-hoc analysis from two complementary perspectives: (1) mask-level comparison and (2) comparison against cross-attention maps.
Mask-level comparison.
Our first goal is to verify that TDM indeed identifies the correct structural region to be edited. Using the crocodile case from Figure 9, we construct a pseudo ground-truth region by taking the union of the source mask and the target mask (generated by SAM [21]), and then downsampling it to the latent resolution. This union mask reflects the intuitive region of change: areas covered by either the original foreground or the edited foreground are exactly those that should be modified, while the background should remain unchanged. We compare this downsampled pseudo mask with our mask $M_{S}$. As shown in the top row of Fig 10, the two maps exhibit similarity in editable areas, providing an intuitive signal that TDM captures the correct semantic region for editing without relying on external supervision. This supports our claim that TDM provides an effective and interpretable estimate of “where” the model intends to apply shape transformation.
TDM vs. cross-attention.
We further compare TDM with cross-attention maps in FLUX DiT. We visualize representative timesteps from the most responsive attention block. Because cross-attention activation is tightly tied to prompt tokens, its localization quality depends heavily on which word has the strongest response. In practice, identifying a single “correct” token is not feasible—responses vary significantly across tokens, heads, and layers—so we use a softmax-normalized variant over all text tokens. Even under this stabilized setting, cross-attention remains spatially noisy and useless with the true editing region. These attention maps are extracted when using TDM-guided region control. Maps without guided control are substantially noisier and therefore omitted for clarity. In contrast, TDM offers a much cleaner and more direct indication of structural change, and it is obtained in a far simpler manner by relying solely on trajectory differences.
Figure 10: Post-hoc analysis of structural localization signals.